# New possibilities in identification of binary choice models with fixed effects

### v5

@@ A.1 Proof of Theorem 1
A.2 Proof of Theorem 2
@@ A.5 Proof of Theorem 5
A.6 Proof of results in Section 5
@@
We prove the first claim since the proof of the second claim is analogous.
The “only if” part is easy. Assume that $\phi(X)=E(Y_{1}-Y_{0}\mid X)\leq 0$ almost surely. Then $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}\phi(X)\leq 0$ for any $q\in\mathbb{R}^{K}$.
We now prove the “if” part. Suppose that $\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq 0$. We consider $q=\beta$. By Lemma 1, $\rho(\beta)=E\mathbf{1}\{W^{\prime}\beta\geq 0\}(Y_{1}-Y_{0})=E\mathbf{1}\{W^{\prime}\beta\geq 0\}\cdot\phi(X)=E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)$. Since $\rho(\beta)\leq\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq 0$, it follows that $P(\phi(X)>0)=0$, which means that $\phi(X)\leq 0$ almost surely. ∎
@@
We now show the validity of the test. Let $c_{1-\alpha}^{*}$ be such that $P\left(\sup_{q\in\mathbb{R}^{K}}S_{*}(q)>c_{1-\alpha}^{*}\right)=\alpha$. If $\tau_{1}\leq-\tau_{2}$, then
@@
Therefore, under the null hypothesis of $\min\{\tau_{1},-\tau_{2}\}\leq 0$, if $\tau_{1}\leq-\tau_{2}$, we have
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}^{*}\right)$
$\leq\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-\min\{\tau_{1},-\tau_{2}\}\right)>c_{1-\alpha}^{*}\right)$
[…]
where (i) follows by the continuous mapping theorem and the fact that $S_{n}(\cdot)$ converges weakly to $S_{*}(\cdot)$.
@@
Therefore, under the null hypothesis of $\min\{\tau_{1},-\tau_{2}\}\leq 0$, if $\tau_{1}>-\tau_{2}$, we have
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}^{*}\right)$
$\leq\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\left(\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}-\min\{\tau_{1},-\tau_{2}\}\right)>c_{1-\alpha}^{*}\right)$
$\leq\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}\left(-S_{n}(q)\right)>c_{1-\alpha}^{*}\right)$
[…]
where (i) follows by the fact that $-S_{*}(\cdot)$ and $S_{*}(\cdot)$ have the same distribution because $S_{*}(\cdot)$ is a mean-zero Gaussian process.
By (27) and (28), no matter whether $\tau_{1}\leq-\tau_{2}$ or $\tau_{1}>-\tau_{2}$, we always have that under the null hypothesis of $\min\{\tau_{1},-\tau_{2}\}\leq 0$,
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}^{*}\right)\leq\alpha.$
This proves the size control since we already proved that $S_{*}(\cdot)$ can be bootstrapped. To see the power, we notice that $\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}=\min\{\tau_{1},-\tau_{2}\}+o_{P}(1)$. Thus, under a fixed alternative of $\min\{\tau_{1},-\tau_{2}\}\geq c_{*}>0$, we would have $\sqrt{n}\min\{\tau_{1},-\tau_{2}\}\geq\sqrt{n}c_{*}\rightarrow\infty$, which implies that
$\limsup_{n\rightarrow\infty}P\left(\sqrt{n}\min\{\hat{\tau}_{1},-\hat{\tau}_{2}\}>c_{1-\alpha}^{*}\right)=1.$

### v7

@@ A.1 Proof of Theorem 1
A.2 Proof of Corollary 1
Proof of Corollary 1.
Let $d=\dim(X_{t})$. We view this case as a special case of Theorem 1 once we treat $\tilde{X}_{t}=(X_{t}^{\prime},\mathbf{1}\{t=0\})^{\prime}$ as the covariate, where the vector of coefficients is $\tilde{\beta}=(\beta^{\prime},0)^{\prime}$. Since the support of $X_{1}-X_{0}$ contains an open neighborhood of zero, we can restrict the observations to those such that the support of $X_{1}-X_{0}$ is an open neighborhood of zero. Moreover, the sign saturation condition holds because the sign of $E(Y_{1}-Y_{0}\mid X)$ is the same as the sign of $(X_{1}-X_{0})^{\prime}\beta$. By Theorem 1, the vector $(\beta^{\prime},0)^{\prime}$ is identified up to scale. Therefore, $\beta$ is identified up to scale. ∎
A.3 Proof of Theorem 2
@@ A.6 Proof of Theorem 5
A.7 Proof of results in Section 4
@@
Notice that $\rho(q)=E\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})=E\mathbf{1}\{W^{\prime}q\geq 0\}\cdot\phi(X)$. Then by construction $\rho(q)\leq E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)$ for any $q$, which means
$\sup_{q\in\mathbb{R}^{K}}\rho(q)\leq E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X).$
On the other hand,
[…]
where (i) follows by Lemma 1. The above two displays imply $\sup_{q\in\mathbb{R}^{K}}\rho(q)=E\mathbf{1}\{\phi(X)\geq 0\}\cdot\phi(X)=E\max\{\phi(X),0\}$. The proof of the other claim follows analogously with $\inf_{q}\rho(q)=\rho(-\beta)=E\min\{\phi(X),0\}$. ∎
@@
Let $c_{1-\alpha}^{*}$ be such that $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}S_{*}(q)>c_{1-\alpha}^{*}\right)=\alpha$. If $\tau_{1}\leq-\tau_{2}$, then
@@
Therefore,
$\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)\leq\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\}.$
Therefore,
$P\left(\sqrt{n}\left(\hat{\tau}_{*}-\tau_{*}\right)>c_{1-\alpha}^{*}\right)\\ \leq P\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)>c_{1-\alpha}^{*}\right)\cdot\mathbf{1}\{\tau_{1}\leq-\tau_{2}\}+P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))>c_{1-\alpha}^{*}\right)\cdot\mathbf{1}\{\tau_{1}>-\tau_{2}\}.$
Since $S_{n}(\cdot)$ converges weakly to $S_{*}(\cdot)$, $P\left(\sup_{q\in\mathbb{R}^{K}}S_{n}(q)>c_{1-\alpha}^{*}\right)$ and $P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{n}(q))>c_{1-\alpha}^{*}\right)$ converge to $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}(S_{*}(q))>c_{1-\alpha}^{*}\right)$ and $\limsup_{n\rightarrow\infty}P\left(\sup_{q\in\mathbb{R}^{K}}(-S_{*}(q))>c_{1-\alpha}^{*}\right)$, respectively. Both limits are equal to $\alpha$; this is because $-S_{*}(\cdot)$ and $S_{*}(\cdot)$ have the same distribution. The result follows since we already proved that $S_{*}(\cdot)$ can be bootstrapped.∎